Circular No. 6320

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

4U 1728–34 T. Strohmayer and W. Zhang, Universities Space Research Associa-

tion (USRA) and the Laboratory for High Energy Astrophysics, Goddard Space Flight Center (LHEA/GSFC); and J. Swank, LHEA/GSFC, report the detection with the proportional counter array (PCA) onboard the X-ray Timing Explorer (XTE) of very-high-frequency variability in the quiescent x-ray intensity from the low-mass x-ray binary 4U 1728–34: “Two features, perhaps distinct, are seen in the power spectrum and appear to be correlated with the source intensity. When the PCA source countrate is 1500–1800 counts/s, a quasiperiodic oscillation (QPO) is seen with a centroid frequency of 800 Hz, FWHM ∼ 94 Hz, and fractional rms variation in the x-ray intensity of 8.4 percent. This feature is accompanied by strong, high-frequency noise extending up to 200 Hz. When the PCA source countrate > 2100 counts/s, a much narrower power spectral feature is seen at 740–800 Hz, with a FWHM of only 8 Hz. The strength of this feature corresponds to a fractional rms variation of ∼ 5 percent. The high-frequency noise seen in the low-countrate state is not present in the high-countrate state. Preliminary searches with data sampled at 16 μs do not reveal any additional significant features in the power spectrum above 1 kHz. Analysis of data segments separated in time suggests that the center frequencies of the features can vary by as much as 5–6 percent over several hours. To our knowledge, this represents the highest-frequency time variability yet seen from a neutron-star system. Explanations in terms of either keplerian frequencies or a beat-frequency model cannot yet be ruled out, although no evidence has yet been seen for a coherent pulsar frequency in the same data. In addition, we note that oscillation modes associated with the crust of a neutron star have frequencies in the range observed.”

COMET 29P/SCHWASSMANN-WACHMANN 1 This comet appears to be in outburst, as indicated by the following to-

tal magnitude estimates (visual unless otherwise noted): Jan. 4.83 UT, 13.8 (Kojima, Chiyoda, Japan, 0.25-m reflector + CCD); 21.34, 13.2 (A. Hale, Cloudcroft, NM, 0.41-m reflector); Feb. 2.64, 14.2 (Kojima); 13.25, 12.6 (R. Keen, Mt. Thorodin, CO, 0.32-m reflector; 2′.5 coma); 15.66, 13.3 (Kojima); 18.06, 13.8 (J. Carvajal, Teruel, Spain, 0.44-m reflector); 19.54, 11.6 (Kojima; strong, nearly stellar condensation with a faint coma extended toward the northwest); 20.27, 11.8 (Hale; 0′.2 coma).

1996 February 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green